Prove Sin(90°+ X) - Sin (90° - X) = 0
Answer by jim_thompson5910(35256)

sin(90° + x) - sin(90° - x) = 0

[sin(90° + x)] - [sin(90° - x)] = 0

[sin(90°)cos(x) + cos(90°)sin(x)] - [sin(90°)cos(x) - cos(90°)sin(x)] = 0

sin(90°)cos(x) + cos(90°)sin(x) - sin(90°)cos(x) + cos(90°)sin(x) = 0

[sin(90°)cos(x) - sin(90°)cos(x)] + [cos(90°)sin(x) + cos(90°)sin(x)] = 0

0*sin(90°)cos(x) + 2*cos(90°)sin(x) = 0

0 + 2*0*sin(x) = 0

0 + 0*sin(x) = 0

0 + 0 = 0

0 = 0